# SciTS: Scientific Time Series Understanding and Generation with LLMs

### v1

@@ Abstract
The scientific reasoning ability of large language models (LLMs) has recently attracted significant attention. Time series, as a fundamental modality in scientific data, presents unique challenges that are often overlooked in current multimodal LLMs, which either encode numerical sequences as text or convert them into images. Such approaches may be insufficient for comprehensive scientific time series understanding and generation. Existing unified time series models typically specialise in either forecasting or analysis, and their effectiveness on non-periodic, heterogeneous scientific signals remains unclear. To address these gaps, we introduce SciTS, a benchmark spanning 12 scientific domains and 43 tasks, with over 50k+ instances, both univariate and multivariate signals ranging from $10^{0}$ to $10^{7}$ in length and up to 10 MHz in frequency. We benchmark 17 models, including text-only LLMs, multimodal LLMs, and unified time series models, and find that general-purpose LLMs exhibit stronger generalisability than specialised time series models, while representing time series as text or images limits their performance due to excessively long sequences and loss of numerical precision, respectively. We then introduce TimeOmni, a framework that equips LLMs with the ability to understand and generate time series while remaining compatible with general-purpose LLM training. This work fills a gap in both dedicated benchmarks and modelling frameworks for scientific time series, paving the way for LLMs to understand and generate complex temporal scientific data. [^1] Code and data will be made available upon acceptance.
@@ 1 Introduction
To address these gaps, we first introduce SciTS, a large-scale benchmark for scientific time series understanding and generation. SciTS comprises 52,056 instances spanning 43 domain-specific tasks across 12 scientific disciplines: Astronomy, Bioacoustics, Earth Science, Economics, Energy, Manufacturing, Mathematics, Meteorology, Neuroscience, Physiology, Radar, and Urbanism. The benchmark covers both univariate and multivariate signals, with lengths ranging from $10^{0}$ to $10^{7}$ and frequencies up to 10 MHz, and includes 7 major task types: anomaly detection, classification, multiple-choice question answering, event localisation, forecasting, imputation, and synthesis. Furthermore, we propose TimeOmni, the first LLM-based framework for unified scientific time series understanding and generation. TimeOmni leverages the reasoning and knowledge of LLMs while explicitly modelling temporal dynamics. Given an input time-series signal and a task prompt, it can generate textual responses or time-series outputs. TimeOmni employs multiple patch experts with a novel routing mechanism that automatically selects the most suitable patch expert, enabling it to handle diverse signals with varying lengths, resolutions, and dimensions. The framework can be readily incorporated into general-purpose LLMs for joint training with other modalities and tasks.
@@
Summary of Contributions. 1) We introduce SciTS, to our knowledge the most comprehensive benchmark to date for scientific time series, spanning a wide range of domains, tasks, and signal types. 2) We present TimeOmni, the first LLM-based framework that explicitly models temporal dynamics and supports both understanding and generation of diverse time-series data. 3) We conduct a large-scale evaluation on SciTS, revealing key challenges in generalising to scientific domains and efficiently handling scientific time series.
@@ 2 Related Work
SciTS TimeSeries 7 $\surd$ 12 Real & Synthetic $\surd$ 52,056
Representation Learning for Time Series. Developing unified time series modelling remains an open challenge. Time series span diverse domains and exhibit varied scales, sampling rates, and dynamic patterns, complicating representation learning and posing challenges for models to adapt across tasks and domains. Transformer-based architectures form the backbone of unified time series models, including encoder-focused designs (Woo et al., 2024), decoder-based structures (Liu et al., 2025; Ansari et al., 2024; Wang et al., 2025b), and their extensions with mixture-of-experts (Xu et al., 2025a; Shi et al., 2025). LLMs have been recently explored for time series analysis, though their effectiveness remains debated: LLM-based forecasters often perform no better than strong non-LLM baselines, highlighting the difficulty of transferring linguistic pretraining to temporal prediction (Tan et al., 2024; Zhou & Yu, 2025). To bridge this gap, multimodal approaches treat time series as an additional modality (Xie et al., 2025), and alignment-based methods adapt temporal signals into language-compatible forms (Hu et al., 2025; Zhang et al., 2025; Jin et al., 2024). Despite these advances, existing approaches largely focus on either understanding or generation, rather than a unified treatment of both. UniTS (Gao et al., 2024) makes a step forward by integrating QA and forecasting tasks, while it relies on a separate architectural design, not compatible with general-purpose LLM training. Moreover, scientific signals, which are often heterogeneous and non-periodic, remain unexplored. In this work, we propose TimeOmni, a framework that unifies 7 types of understanding and generation tasks, and can be readily integrated into LLMs for joint training with other modalities and tasks.
@@ 3 SciTS Dataset and Tasks
The SciTS benchmark consists of 52,056 samples spanning 43 tasks across 12 disciplines: Astronomy, Bioacoustics, Earth Science, Economics, Energy, Manufacturing, Mathematics, Meteorology, Neuroscience, Physiology, Radar, and Urbanism, as shown in Fig. 2. Time series from different domains exhibit widely varying characteristics: frequencies range from once per day in light curve observations to kilohertz for bird chirps in bioacoustics, and even megahertz in radar communications; lengths vary from fewer than 10 points for gait-freezing measurements to millions of points for animal sounds; moreover, SciTS includes multivariate signals, with the number of channels reaching up to 58 for electroencephalogram (EEG) recordings in neuroscience. Detailed dataset statistics are presented in Table 2, and more details can be found in Appendix D.2.
@@ 3.1 Data Collection
# Instances 52,056
    # Understanding tasks 28,677
    # Classification 17,964
@@
    # Anomaly detection 9,997
   # Generation tasks 23,379
@@ 3.3 Type of Tasks
4 TimeOmni for Time Series Understanding and Generation
@@
The structure of TimeOmni is shown in Fig. 3, consisting of a Time Series Encoder, a LLM backbone, and output heads for diverse signals. Given an input time series signal $\mathbf{X}\in\mathbb{R}^{T^{\prime}\times N}$, it is first flattened along the time dimension to obtain $\mathbf{X}^{\prime}\in\mathbb{R}^{NT^{\prime}\times 1}$. The encoder automatically selects a patch expert based on $NT^{\prime}$, producing $\mathbf{X}_{\text{enc}}\in\mathbb{R}^{T_{\text{enc}}\times D_{\text{llm}}}$, where $D_{\text{llm}}$ is the LLM hidden dimension and $T_{\text{enc}}$ is typically 100 to 200. The task prompt is tokenised and embedded into $\mathbf{P}\in\mathbb{R}^{L\times D_{\text{llm}}}$, which is concatenated with $\mathbf{X}_{\text{enc}}$ as input to the LLM backbone. Output embeddings from the LLM are processed according to task type: a softmax layer to generate discrete text tokens for understanding tasks and a regression head to output a time series for generation tasks.
@@ 4.1 Time Series Encoder
As shown in Fig. 3, the Time Series Encoder consists of three main components: a Router, a Patch Expert family, and a Patch Reprogramming module (Jin et al., 2024). For a flattened input of length $T=NT^{\prime}$, the Router selects a patch size $D_{\text{patch}}$ to ensure that the output sequence length after the Patch Expert layer falls between 100 and 200, i.e., ${T}/{200}<D_{\text{patch}}<{T}/{100}$. The Patch Expert first patchifies the signal, reshaping $\mathbf{X}^{\prime}$ from $\mathbb{R}^{T\times 1}$ to $\mathbb{R}^{\lceil T/D_{\text{patch}}\rceil\times D_{\text{patch}}}$. A 1-dimensional (-dim) convolution then maps it to $\mathbf{X}_{\text{patch}}\in\mathbb{R}^{\lceil T/D_{\text{patch}}\rceil\times D_{\text{enc}}}$, where $D_{\text{enc}}$ is consistent across the Patch Expert family. The Patch Reprogramming module re-represents the time series using the LLM’s vocabulary embeddings $\mathbf{E}\in\mathbb{R}^{\text{vocab\_size}\times D_{\text{llm}}}$. $\mathbf{E}$ is first projected through a linear layer to $\mathbb{R}^{1000\times D_{\text{llm}}}$, then $\mathbf{X}_{\text{patch}}$ attends to $\mathbf{E}$ via multi-head cross-attention, with $\mathbf{X}_{\text{patch}}$ as query and $\mathbf{E}$ as key and value. A final linear projection produces the encoder output $\mathbf{X}_{\mathrm{enc}}$.
@@ 4.2 LLM Input and Output
For understanding tasks, we use a Prompt-as-suffix strategy, where $\mathbf{X}_{\text{enc}}$ is followed by $\mathbf{P}$. For generation tasks, a Prompt-as-prefix approach is adopted, placing $\mathbf{P}$ before $\mathbf{X}_{\text{enc}}$. The concatenated vector is fed into the LLM. For understanding tasks, output embeddings of LLMs pass through a softmax layer to produce discrete text tokens, following the standard LLM generation process. For generation tasks, the output embeddings are flattened and mapped via a linear layer to the required length of the output time series. To handle the wide range of output lengths in the benchmark, we predefine a set of regression heads and the model selects the one with the closest length, truncating the output if necessary.
@@ 6.1 Task Coverage and Success Rate
On the other hand, unified time series models typically focus on a certain type of task, resulting in low overall task coverage (Fig. 4(a)). For example, Moirai and TimeMoE only support forecasting and cannot handle imputation tasks or tasks requiring text input. Despite this limitation, they achieve high success rates on the tasks they do support, demonstrating the effectiveness of specialised models. Our proposed TimeOmni model, in comparison, successfully handles all tasks and instances, achieving both full coverage and full success across the benchmark.
@@ 6.2 Understanding Tasks
Table 3 compares text-only LLMs, multimodal LLMs, and time series models on understanding tasks across different scientific disciplines, with results reported as the average F1 score across tasks in each discipline. Overall, models achieve limited performance on scientific domains such as astronomy, neuroscience, and physiology, possibly due to these scientific domains being largely absent from the training data of general-purpose models. Performance is particularly low for bioacoustics and radar, with F1 scores below 10%, highlighting the challenging nature of the SciTS benchmark. In contrast, nearly all models achieve strong performance on economics (except LLaMA3), as this domain is closer to general knowledge and thus falls within the scope of general-purpose models. It is worth noting that for the urbanism domain, the results in Table 3 are primarily from Task URU04222The mapping of task IDs to their corresponding tasks is provided in Table 6., an anomaly detection task with relatively few instances. Most models effectively perform random guessing, and the highest F1 score (bolded) largely reflects a bias toward a single class. In fact, GPT-4.1-mini and Qwen-2.5-VL achieve higher accuracy with detailed results provided in Appendix G.
Overall, multimodal LLMs that take images as input outperform text-only LLMs, possibly because tasks requiring high-level understanding do not strongly depend on precise numerical values, and representing time series as text can produce extremely long inputs that are difficult for LLMs to process. Converting time series into images compresses these long sequences more effectively. We also found models specialised for time series do not necessarily surpass general-purpose LLMs across many domains, which indicates that when exposed to data domains unseen during training, LLMs demonstrate stronger generalisation capabilities. Across most domains, the proposed TimeOmni model ranks first, achieving the overall best performance, with Gemini2.5-Flash ranking second.
@@ 6.3 Generation Tasks
Table 4 compares various models on generation tasks. Overall, generation tasks prove more challenging, as reflected by the large swMAPE values reported in the table. Unlike understanding tasks, representing time series as text generally outperforms image-based inputs for generation tasks where precise numerical values are critical for accurate generation, as expected. Gemini2.5 performs best in neuroscience, physiology, and urbanism, where the output time series are relatively short (typically fewer than 20 steps [^3] Detailed statistics of each task can be found in Table 7.). For long-term forecasting tasks, such as Task ENG02 with output lengths up to 720, all LLMs struggle to generate sequences of the required length, with success rates commonly below 10% (Table 12). Closed-source LLMs continue to outperform open-source counterparts across most domains. Many values for unified forecasting models are missing in the table, as they cannot handle imputation or event localisation tasks. However, they achieve superior performance on supported forecasting tasks, with Moirai giving the overall lowest average swMAPE in economics and mathematics. The proposed TimeOmni model achieves the highest overall ranking, closely followed by Gemini2.5-Flash using text input. Detailed results for each task are provided in Appendix G.
@@ 6.4 Model Rankings Across Disciplines
Fig. 5 shows the average rankings of 17 tested models across 12 disciplines. We rank models for each of the 43 tasks and average the ranking over tasks within each discipline. The benchmarking results reveal several clear patterns. First, the proposed TimeOmni model achieves consistently top performance across nearly all disciplines, with rankings between first and second in most domains and only slightly lower in mathematics and urbanism. General-purpose LLMs such as GPT-4.1-mini, Gemini2.5-Flash, DeepSeek-V3, and GPT-5-mini deliver strong all-around results, usually falling within the top 5 to 8 ranks. Multimodal LLMs like InternVL and QwenVL occupy a middle ground, showing competitive performance in sensory-heavy fields such as bioacoustics, earth science, and radar, but weaker results in other areas. Specialised time-series models, including Moirai, TimeMoE, Chronos, and ChatTS, struggle with generalisation. They often collapse to the bottom of the ranking across most domains due to limitations of supported task types, but occasionally excel in specific cases, such as Moirai and Chronos in mathematics and ChatTS in earth science. Discipline-specific patterns also emerge: models consistently perform well across language-heavy domains such as Economics, while dedicated time-series models show strengths in more structured or periodic domains like Energy and Meteorology, but degrade sharply elsewhere.
@@
These results suggest two key insights. First, time series should be treated as a distinct modality rather than being directly converted to text or images, as such transformations lead to information loss and limit model performance. Second, general-purpose LLMs demonstrate better generalisation to unseen domains compared with specialised time-series models, highlighting the benefit of combining LLMs with dedicated time series processing to effectively handle diverse scientific data. The proposed TimeOmni addresses this gap by providing a framework that incorporates time series into LLMs while remaining compatible with general LLM training.

### v2

@@ Abstract
The scientific reasoning ability of large language models (LLMs) has recently attracted significant attention. Time series, as a fundamental modality in scientific data, presents unique challenges that are often overlooked in current multimodal LLMs, which either encode numerical sequences as text or convert them into images. Such approaches may be insufficient for comprehensive scientific time series understanding and generation. Existing unified time series models typically specialise in either forecasting or analysis, and their effectiveness on non-periodic, heterogeneous scientific signals remains unclear. To address these gaps, we introduce SciTS, a benchmark spanning 12 scientific domains and 43 tasks, with over 50k+ instances, both univariate and multivariate signals ranging from $10^{0}$ to $10^{7}$ in length and up to 10 MHz in frequency. We benchmark 17 models, including text-only LLMs, multimodal LLMs, and unified time series models, and find that general-purpose LLMs exhibit stronger generalisability than specialised time series models, while representing time series as text or images limits their performance due to excessively long sequences and loss of numerical precision, respectively. We then introduce TimeOmni, a working example to explore insights into how LLMs can be extended to handle scientific time series while remaining compatible with general-purpose LLM training. This work fills a gap in both dedicated benchmarks and illustrative frameworks for scientific time series, paving the way for LLMs to understand and generate complex temporal scientific data. [^1] SciTS benchmark is available at: https://huggingface.co/datasets/OpenTSLab/SciTS. TimeOmni framework is available at: https://github.com/OpenTSLab/TimeOmni.
@@ 1 Introduction
To address these gaps, we first introduce SciTS, a large-scale benchmark for scientific time series understanding and generation. SciTS comprises 54,023 instances spanning 43 domain-specific tasks across 12 scientific disciplines: Astronomy, Bioacoustics, Earth Science, Economics, Energy, Manufacturing, Mathematics, Meteorology, Neuroscience, Physiology, Radar, and Urbanism. The benchmark covers both univariate and multivariate signals, with lengths ranging from $10^{0}$ to $10^{7}$ and frequencies up to 10 MHz, and includes 7 major task types: anomaly detection, classification, multiple-choice question answering, event localisation, forecasting, imputation, and synthesis.
@@
In addition, we introduce TimeOmni as a working example to explore the key ingredients that may be necessary for unified scientific time series understanding and generation. TimeOmni leverages the reasoning and knowledge of LLMs while explicitly modelling temporal dynamics. Given an input time-series signal and a task prompt, it generates textual responses or time-series outputs. In order to handle diverse signals with varying lengths, resolutions, and dimensions, TimeOmni employs multiple patch experts with a novel routing mechanism that automatically selects the most suitable patch expert. The framework is intended to be readily incorporated into general-purpose LLMs for joint training with other modalities and tasks.
Summary of Contributions. 1) We introduce SciTS, to our knowledge the most comprehensive benchmark to date for scientific time series, spanning a wide range of domains, tasks, and signal types. 2) We conduct a large-scale evaluation on SciTS, revealing key challenges in generalising to scientific domains and efficiently handling scientific time series. 3) We introduce TimeOmni to study what is required for LLMs to handle diverse time-series data, revealing insights into explicit and adaptive temporal modelling.
@@ 2 Related Work
SciTS TimeSeries 7 $\surd$ 12 Real & Synthetic $\surd$ 54,023
Representation Learning for Time Series. Developing unified time series modelling remains an open challenge. Time series span diverse domains and exhibit varied scales, sampling rates, and dynamic patterns, complicating representation learning and posing challenges for models to adapt across tasks and domains. Transformer-based architectures form the backbone of unified time series models, including encoder-focused designs (Woo et al., 2024), decoder-based structures (Liu et al., 2025; Ansari et al., 2024; Wang et al., 2025b), and their extensions with mixture-of-experts (Xu et al., 2025a; Shi et al., 2025). LLMs have been recently explored for time series analysis, though their effectiveness remains debated: LLM-based forecasters often perform no better than strong non-LLM baselines, highlighting the difficulty of transferring linguistic pretraining to temporal prediction (Tan et al., 2024; Zhou & Yu, 2025). To bridge this gap, multimodal approaches treat time series as an additional modality (Xie et al., 2025), and alignment-based methods adapt temporal signals into language-compatible forms (Hu et al., 2025; Zhang et al., 2025; Jin et al., 2024). Despite these advances, existing approaches largely focus on either understanding or generation, rather than a unified treatment of both. UniTS (Gao et al., 2024) makes a step forward by integrating QA and forecasting tasks, while it relies on a separate architectural design, not compatible with general-purpose LLM training. Moreover, scientific signals, which are often heterogeneous and non-periodic, remain unexplored. In this work, we propose TimeOmni, a framework that unifies 7 types of understanding and generation tasks, and can be readily integrated into LLMs for joint training with other modalities and tasks. Extended discussion on the related work is provided in Appendix B.
@@ 3 SciTS Dataset and Tasks
The SciTS benchmark consists of 54,023 samples spanning 43 tasks across 12 disciplines: Astronomy, Bioacoustics, Earth Science, Economics, Energy, Manufacturing, Mathematics, Meteorology, Neuroscience, Physiology, Radar, and Urbanism, as shown in Fig. 2. Time series from different domains exhibit widely varying characteristics: frequencies range from once per day in light curve observations to kilohertz for bird chirps in bioacoustics, and even megahertz in radar communications; lengths vary from fewer than 10 points for gait-freezing measurements to millions of points for animal sounds; moreover, SciTS includes multivariate signals, with the number of channels reaching up to 58 for electroencephalogram (EEG) recordings in neuroscience. Detailed dataset statistics are presented in Table 2, and more details can be found in Appendix D.2.
@@ 3.1 Data Collection
# Instances 54,023
    # Understanding tasks 32,149
    # Classification 18,364
@@
    # Anomaly detection 13,069
   # Generation tasks 21,874
@@ 3.3 Type of Tasks
4 TimeOmni: A Working Example
@@
To explore the key ingredients that may be necessary for extending current LLMs to handle scientific time series, we develop TimeOmni as an illustrative framework. The structure of TimeOmni is shown in Fig. 3, consisting of a time series encoder, a LLM backbone, and output heads for diverse signals. Given an input time series signal $\mathbf{X}\in\mathbb{R}^{T^{\prime}\times N}$, it is first flattened along the time dimension to obtain $\mathbf{X}^{\prime}\in\mathbb{R}^{NT^{\prime}\times 1}$. The encoder automatically selects a patch expert based on $NT^{\prime}$, producing $\mathbf{X}_{\text{enc}}\in\mathbb{R}^{T_{\text{enc}}\times D_{\text{llm}}}$, where $D_{\text{llm}}$ is the LLM hidden dimension and $T_{\text{enc}}$ is typically 100 to 200. The task prompt is tokenised and embedded into $\mathbf{P}\in\mathbb{R}^{L\times D_{\text{llm}}}$, which is concatenated with $\mathbf{X}_{\text{enc}}$ as input to the LLM backbone. Output embeddings from the LLM are processed according to task type: a softmax layer to generate discrete text tokens for understanding tasks and a regression head to output a time series for generation tasks.
@@ 4.1 Time Series Encoder
As shown in Fig. 3, , time series is explicitly modelled by an encoder, which consists of three main components: a router, a patch expert family, and a patch reprogramming module (Jin et al., 2024). For a flattened input of length $T=NT^{\prime}$, the Router selects a patch size $D_{\text{patch}}$ to ensure that the output sequence length after the Patch Expert layer falls between 100 and 200, i.e., ${T}/{200}<D_{\text{patch}}<{T}/{100}$. The Patch Expert first patchifies the signal, reshaping $\mathbf{X}^{\prime}$ from $\mathbb{R}^{T\times 1}$ to $\mathbb{R}^{\lceil T/D_{\text{patch}}\rceil\times D_{\text{patch}}}$. A 1-dimensional (-dim) convolution then maps it to $\mathbf{X}_{\text{patch}}\in\mathbb{R}^{\lceil T/D_{\text{patch}}\rceil\times D_{\text{enc}}}$, where $D_{\text{enc}}$ is consistent across the Patch Expert family. The Patch Reprogramming module re-represents the time series using the LLM’s vocabulary embeddings $\mathbf{E}\in\mathbb{R}^{\text{vocab\_size}\times D_{\text{llm}}}$. $\mathbf{E}$ is first projected through a linear layer to $\mathbb{R}^{1000\times D_{\text{llm}}}$, then $\mathbf{X}_{\text{patch}}$ attends to $\mathbf{E}$ via multi-head cross-attention, with $\mathbf{X}_{\text{patch}}$ as query and $\mathbf{E}$ as key and value. A final linear projection produces the encoder output $\mathbf{X}_{\mathrm{enc}}$.
@@ 4.2 LLM Input and Output
For understanding tasks, we use a Prompt-as-suffix strategy, where $\mathbf{X}_{\text{enc}}$ is followed by $\mathbf{P}$. For generation tasks, a Prompt-as-prefix approach is adopted, placing $\mathbf{P}$ before $\mathbf{X}_{\text{enc}}$. The concatenated vector is fed into the LLM. For understanding tasks, output embeddings of LLMs pass through a softmax layer to produce discrete text tokens, following the standard LLM generation process. For generation tasks, the output embeddings are flattened and mapped via a linear layer to the required length of the output time series. To handle the wide range of output lengths in the benchmark, we predefine a set of regression heads and the model selects the one with the closest length, truncation applied if necessary.
@@ 6.1 Task Coverage and Success Rate
On the other hand, unified time series models typically focus on a certain type of task, resulting in low overall task coverage (Fig. 4(a)). For example, Moirai and TimeMoE only support forecasting and cannot handle imputation tasks or tasks requiring text input. Despite this limitation, they achieve high success rates on the tasks they do support, demonstrating the effectiveness of specialised models. The illustrated TimeOmni framework, in comparison, successfully handles all tasks and instances, achieving both full coverage and full success across the benchmark.
@@ 6.2 Understanding Tasks
Table 3 compares text-only LLMs, multimodal LLMs, and time series models on understanding tasks across different scientific disciplines, with results reported as the average F1 score across tasks in each discipline. Overall, models achieve limited performance on scientific domains such as astronomy, neuroscience, and physiology, possibly due to these scientific domains being largely absent from the training data of general-purpose models. Performance is particularly low for bioacoustics and radar, with F1 scores below 10%, highlighting the challenging nature of the SciTS benchmark. In contrast, nearly all models achieve strong performance on economics (except LLaMA3), as this domain is closer to general knowledge and thus falls within the scope of general-purpose models. It is worth noting that for the urbanism domain, the results in Table 3 are primarily from Task URU04222The mapping of task IDs to their corresponding tasks is provided in Table 6., an anomaly detection task with relatively few instances. Most models effectively perform random guessing, and the highest F1 score (bolded) largely reflects a bias toward a single class. In fact, GPT-4.1-mini and Qwen-2.5-VL achieve higher accuracy with detailed results provided in Appendix F.
Overall, multimodal LLMs that take images as input outperform text-only LLMs, possibly because tasks requiring high-level understanding do not strongly depend on precise numerical values, and representing time series as text can produce extremely long inputs that are difficult for LLMs to process. Converting time series into images compresses these long sequences more effectively. We also found models specialised for time series do not necessarily surpass general-purpose LLMs across many domains, which indicates that when exposed to data domains unseen during training, LLMs demonstrate stronger generalisation capabilities. Across most domains, the illustrated TimeOmni framework achieves the overall best performance, with Gemini2.5-Flash ranking second.
@@ 6.3 Generation Tasks
Table 4 compares various models on generation tasks. Overall, generation tasks prove more challenging, as reflected by the large swMAPE values reported in the table. Unlike understanding tasks, representing time series as text generally outperforms image-based inputs for generation tasks where precise numerical values are critical for accurate generation, as expected. Gemini2.5 performs best in neuroscience, physiology, and urbanism, where the output time series are relatively short (typically fewer than 20 steps [^3] Detailed statistics of each task can be found in Table 7.). For long-term forecasting tasks, such as Task ENG02 with output lengths up to 720, all LLMs struggle to generate sequences of the required length, with success rates commonly below 10% (Table 12). Closed-source LLMs continue to outperform open-source counterparts across most domains. Many values for unified forecasting models are missing in the table, as they cannot handle imputation or event localisation tasks. However, they achieve superior performance on supported forecasting tasks, with Moirai giving the overall lowest average swMAPE in economics and mathematics. The proposed TimeOmni model achieves the highest overall ranking, closely followed by Gemini2.5-Flash using text input. Detailed results for each task are provided in Appendix F.
@@ 6.4 Model Rankings Across Disciplines
Fig. 5 shows the average rankings of 17 tested models across 12 disciplines. We rank models for each of the 43 tasks and average the ranking over tasks within each discipline. The benchmarking results reveal several clear patterns. General-purpose LLMs such as GPT-4.1-mini, Gemini2.5-Flash, DeepSeek-V3, and GPT-5-mini deliver strong all-around results, usually falling within the top 5 to 8 ranks. Multimodal LLMs like InternVL and QwenVL occupy a middle ground, showing competitive performance in sensory-heavy fields such as bioacoustics, earth science, and radar, but weaker results in other areas. Specialised time-series models, including Moirai, TimeMoE, Chronos, and ChatTS, struggle with generalisation. They often collapse to the bottom of the ranking across most domains due to limitations of supported task types, but occasionally excel in specific cases, such as Moirai and Chronos in mathematics and ChatTS in earth science. Discipline-specific patterns also emerge: models consistently perform well across language-heavy domains such as Economics, while dedicated time-series models show strengths in more structured or periodic domains like Energy and Meteorology, but degrade sharply elsewhere. As an illustrative framework, TimeOmni achieves consistently top performance across nearly all disciplines, with slightly lower in mathematics and urbanism.
@@
These results suggest two key insights. First, time series should be treated as a distinct modality rather than being directly converted to text or images, as such transformations lead to information loss and limit model performance. Second, general-purpose LLMs demonstrate better generalisation to unseen domains compared with specialised time-series models, highlighting the benefit of combining LLMs with dedicated time series processing to effectively handle diverse scientific data. TimeOmni provides an illustrative step that incorporates time series into LLMs while remaining compatible with general LLM training.
6.5 Design Analysis: What is Needed for Scientific Time Series Modelling?
To better understand what is essential for modelling scientific time series, we first conducted ablation experiments on the reprogramming module and patch expert routing mechanism of TimeOmni. Due to space constraints, we summarise main findings here and defer complete experimental results to Appendix G. Replacing the reprogramming module with a simple MLP consistently degrades performance, suggesting that explicit reprogramming is important for aligning time series representations with the LLM and improving robustness across tasks. Replacing patch expert routing with a fixed patch size faces a fundamental dilemma: small patch sizes become infeasible for very long sequences due to memory constraints and difficulty of modelling extremely long-range dependencies, while large patch sizes collapse short sequences into a single patch. Choosing an intermediate patch size leads to clear performance drops on sequences with extreme temporal scales, highlighting the need for scale-adaptive patching in scientific time series.
To complement the above analysis, we further investigate whether existing models can be directly adapted to scientific time series through finetuning. We additionally finetuned one image–text LLM (Qwen2.5VL-7B) and one time-series foundation model (TimeMoE) using the same training data as TimeOmni. Finetuning Qwen2.5VL improves instruction-following ability on some tasks, but the gains are limited possibly because representing time series as images compromises numerical precision. We also observed instability during finetuning and difficulty in maintaining balanced performance across tasks, potentially reflecting the limited capacity of image-based representations for time-series data. We then finetune TimeMoE, a forecasting-oriented model. While performance improves on some forecasting tasks, finetuning does not enable it to handle tasks it could not previously address, such as imputation or QA. In sum, results provide evidence that challenges arises from architectural limitations and cannot be overcome through finetuning alone. Full results and discussion are provided in Appendix H.